# ReX: Causal Discovery based on Machine Learning and Explainability techniques

### v1

@@ Appendix B SHAP discrepancy example
Appendix C Performance comparison between single and combinations of regressors
ReX employs two regressors—DFN and GBT—to obtain individual causal graphs from the data (see Section 3.1). The final DAG is constructed by combining these graphs using a union ($\cup$) operation. This operation effectively includes all edges identified by either regressor, ensuring that any edge detected by at least one of them is incorporated into the final solution. This approach aims to leverage the strengths of both regressors, capturing edges that might be missed by one but detected by the other.
Experiments on the synthetic datasets proposed in this study (see Section 4) indicate that this combined approach results in better predictive performance than using the output of a single regressor (Fig. 9). Specifically, the union of the regressors’ outputs yields higher F1 scores across datasets with 10, 15, 20 and 25 variables, suggesting a more accurate recovery of the true causal structures. DFN’s F1 is $0.577$ and GBT’s F1 is $0.59$, while union’s F1 is $0.712$, a $23.4\%$ higher.
@@ Appendix C Performance comparison between single and combinations of regressors
In addition to the union operation, we have also tested the intersection ($\cap$) operation for combining the regressors’ outputs. This method includes only the edges that are identified by both regressors. While the intersection approach also provides better results than using either regressor individually, its performance is slightly inferior ($0.667$) to that of the union operation ($4.9\%$ lower). The more conservative nature of the intersection may exclude some true causal edges identified by only one regressor, leading to a less comprehensive causal graph.
Appendix D Sample DAGs
@@
Fig. 10 and 11 show sample DAGs generated from a Gaussian additive and Sigmoid additive processes, respectively, with $p=10$ variables. Solid green edges represent correctly predicted causal relationships that match the ground truth, light gray dotted lines indicate edges that were present in the ground truth but not predicted by the method, dashed orange lines represent edges where the causal pair was correctly identified but with the wrong direction, and red dash-dot lines correspond to incorrectly predicted edges.
@@
Appendix E Additional results
@@
Figure 12: F1 (a), Precision (b), Recall (c), SHD (d) and SID (e) for ReX as a function of the number of features and synthetic data generation mechanism. Panel (f) represents the training time for the different families of synthetic data and number of variables used (x-axis).

### v2

@@ Appendix C SHAP discrepancy example
Appendix D Performance comparison between single and combinations of regressors
ReX employs two complementary regressors—Deep Feed-forward Networks (DFN) and Gradient Boosting Trees (GBT)—to generate individual causal graph hypotheses from the data (see Section 3.1). The final DAG, $G_{\text{R{e}X}}$, is then constructed by taking the union ($\cup$) of the DAGs derived from each regressor, $G_{DFN}$ and $G_{GBT}$.
The rationale for choosing the union operation is to maximize the discovery of true causal relationships by leveraging the distinct modeling capabilities of DFNs (effective at capturing complex, non-linear functions) and GBTs (robust for tabular data and possessing different inductive biases). Since either type of regressor might uniquely identify certain true edges that the other misses due to its specific learning mechanism, the union aims to create a more comprehensive initial graph, prioritizing the recall of potentially true edges at this stage of the pipeline. It is important to note that each individual graph, $G_{DFN}$ and $G_{GBT}$, is already the result of a robust bootstrapping process (Algorithm 1) designed to stabilize edge selection and reduce spurious connections before the union.
We also evaluated the alternative of using the intersection ($\cap$) of $G_{DFN}$ and $G_{GBT}$. The intersection is a more conservative approach, as it retains only those edges agreed upon by both regressors. This can lead to higher precision but may also discard true causal edges identified by only one model type, potentially resulting in a less comprehensive causal graph.
Experiments on the synthetic datasets proposed in this study (see Section 4) indicate that the union approach results in better overall performance. As shown in Fig. 11, the union of the regressors’ outputs yields higher F1 scores (0.712) across datasets with 10, 15, 20, and 25 variables compared to both individual regressors (DFN F1: $0.577$; GBT F1: $0.59$) and the intersection strategy (F1: $0.667$). Specifically, the union’s F1 score is $23.4\%$ higher than the average F1 score of the individual DFN and GBT models, and $4.9\%$ higher than that of the intersection approach. This suggests that, for the datasets and configurations tested, the union provides a superior balance of precision and recall, capturing more true causal structures overall.
@@ Appendix D Performance comparison between single and combinations of regressors
While the union carries a theoretical risk of including more spurious edges than the intersection if one model is prone to specific errors not fully mitigated by its individual bootstrapping phase, the subsequent cycle removal and edge orientation steps (Section 3.3 and Section 3.4), particularly the SHAP discrepancy measure for resolving conflicts, further refine the graph. The empirically superior F1-score ultimately guided our decision to adopt the union approach as the default strategy in ReX.
Appendix E Sample DAGs
@@
Fig. 12 and 13 show sample DAGs generated from a Gaussian additive and Sigmoid additive processes, respectively, with $p=10$ variables. Solid green edges represent correctly predicted causal relationships that match the ground truth, light gray dotted lines indicate edges that were present in the ground truth but not predicted by the method, dashed orange lines represent edges where the causal pair was correctly identified but with the wrong direction, and red dash-dot lines correspond to incorrectly predicted edges.
@@
Appendix F Additional results
@@
Figure 14: F1 (a), Precision (b), Recall (c), SHD (d) and SID (e) for ReX as a function of the number of features and synthetic data generation mechanism (Appendix G). Panel (f) represents the training time for the different families of synthetic data and number of variables used (x-axis).
Appendix G Synthetic data generation
Each synthetic dataset is built with $p$ variables ($p=10,15,20,25$) and $m=500$ samples.
  1. 1.
    The DAG structure is such that the number of parents for each variable is uniformly drawn in $\{0,\ldots,5\}$;
  2. 2.
    For the $i$-th DAG, the mean $\mu_{i}$ and variance $\sigma_{i}$ of the noise variables are drawn as $\mu_{i}\sim\mathcal{U}(-2,2)$ and $\sigma_{i}\sim\mathcal{U}(0,0.4)$, and the distribution of the noise variables is set to $\mathcal{N}(\mu_{i},\sigma_{i})$;
  3. 3.
    For each graph, a $500$-sample dataset is i.i.d. generated following the topological order of the graph, with, for $\ell=1,\ldots,500$,
    $x^{(\ell)}=\bigl(x_{1}^{(\ell)},\ldots,x_{p}^{(\ell)}\bigr),\qquad x_{i}^{(\ell)}\sim f_{i}\bigl(X_{\operatorname{Pa}(i)},\varepsilon_{i}\bigr),\qquad E_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}).$
All variables are normalized to zero mean and unit variance.
Five categories of causal mechanisms have been considered, replicating those described in [26]:
  1. I.
    Linear: $X_{i}=\sum_{j\in\operatorname{Pa}(i)}a_{i,j}\,X_{j}+\varepsilon_{i}$,   where $a_{i,j}\sim\mathcal{N}(0,1)$.
  2. II.
    Polynomial: $X_{i}=\sum_{j\in\operatorname{Pa}(i)}c_{i,j}\,X_{j}^{d}+\varepsilon_{i}$,   where $c_{i,j}\sim\mathcal{N}(0,1)$, and $d$ is the degree of the polynomial.
  3. III.
    GP AM: $X_{i}=\sum_{j\in\operatorname{Pa}(i)}f_{i,j}(X_{j})+\varepsilon_{i}$,   where $f_{i,j}$ is a univariate Gaussian process with a Gaussian kernel of unit bandwidth.
  4. IV.
    GP Mix: $X_{i}=f_{i}\!\bigl([X_{\operatorname{Pa}(i)},\,\varepsilon_{i}]\bigr)$,   where $f_{i}$ is a multivariate Gaussian process with a Gaussian kernel of unit bandwidth.
  5. V.
    Sigmoid Mix: $X_{i}=f_{i}\Bigl(\displaystyle\sum_{j\in\operatorname{Pa}(i)}X_{j}\Bigr)+\varepsilon_{i}$, where $f_{i}$ is:
    $f_{i,j}(x_{j})=a\,\frac{b\,(x_{j}+c)}{1+b\,(x_{j}+c)},\qquad a\sim\operatorname{Exp}(4)+1,\;b\sim\mathcal{U}\!\bigl([-2,-0.5]\cup[0.5,2]\bigr),\;c\sim\mathcal{U}([-2,2]).$
Appendix H Computational complexity
The computational cost of ReX, as depicted in Figure 1 (overall workflow) and detailed in Algorithms 1 and 2, is influenced by several stages. Understanding these helps contextualize the empirical runtime results presented in Figure 3d.
The primary computational stages of ReX are:
  • Model Training (Section 3.1): This involves training $p$ predictive models for each of the two regressor types (DFN and GBT), where $p$ is the number of variables. The cost is $O(p\cdot(C_{\text{HPO}}+C_{\text{train}}))$, where $C_{\text{HPO}}$ is the cost of hyperparameter optimization and $C_{\text{train}}$ is the cost of training a single regressor (DFN or GBT) on $m$ samples and p1p-1 features. This depends on the chosen model’s complexity (e.g., network architecture, number of trees) and training epochs.
  • Bootstrapped SHAP Value Computation (Algorithm 1): This stage iterates $T$ times. In each iteration, for every one of the $p$ target variables, SHAP values are computed for the p1p-1 input features using a subset of $m^{\prime}$ samples. The cost of computing SHAP values, $C_{\text{SHAP}}$, varies: TreeExplainer for GBTs is generally efficient, scaling with $m^{\prime}$, $p$, number of trees, and tree depth (e.g., approximately $O(m^{\prime}\cdot N_{trees}\cdot D\cdot p)$ for $m^{\prime}$ samples explained over $p$ features). GradientExplainer for DFNs can be more intensive, often scaling with $m^{\prime}$, $p$, the number of samples for expectation, and the cost of model evaluation. This makes the overall SHAP computation roughly $O(T\cdot p\cdot C_{\text{SHAP}})$.
  • Parent Selection (Algorithm 2): For each target variable within each bootstrap iteration, this algorithm processes p1p-1 SHAP values. It involves pairwise distance calculations and an iterative DBSCAN clustering approach. Given its current structure involving iterative adjustments of $\zeta$ based on distances, its complexity is estimated to be polynomial in $p$, potentially around $O(p^{3})$.
  • Edge Orientation (Section 3.3): This is applied to $E_{\text{undir}}$ candidate edges (where $E_{\text{undir}}\leq p(p-1)/2$). It involves fitting two regression models (GAMs are employed for efficiency with $m$ samples) and performing two HSIC tests per edge. HSIC tests typically scale as $O(m^{2})$ (or $O(m\log m)$ with approximations). The overall complexity for this stage is roughly $O(E_{\text{undir}}\cdot(C_{\text{GAM}}+C_{\text{HSIC}}))$.
  • Final DAG Construction (Section 3.4): This includes graph union operations ($O(p^{2})$) and cycle detection/resolution. The cost of resolving cycles using SHAP discrepancy depends on the number and complexity of cycles and whether per-sample SHAP values ($\phi_{k,j}$ in Equation 12) are recomputed or readily available.
The dominant computational burden in ReX typically arises from the bootstrapped SHAP value computation, especially as $p$ and $m$ increase, leading to a high-degree polynomial scaling in $p$. This theoretical expectation aligns with the empirical runtime trends observed in Figure 3d. Compared to other methods:
  • Constraint-based methods (PC [2], FCI [12]): These can be exponential in $p$ in the worst-case due to the need to test conditional independencies over many subsets, though they are often faster for sparse graphs. ReX avoids this exhaustive CI testing by leveraging feature importance from global predictive models.
  • Score-based methods (GES [13]): These perform a heuristic search in the space of DAGs, and their complexity depends on the size of this space and the cost of evaluating the score for each candidate structure.
  • SCM-based methods: LiNGAM [14] is relatively efficient (e.g., $O(p^{3}+p^{2}m)$) under its specific assumptions (linear, non-Gaussian). CAM [15] can be computationally intensive due to its search for parent sets and fitting of non-linear additive models.
  • Continuous optimization methods (NOTEARS [17]): These typically have polynomial complexity (e.g., $O(mp^{2}+p^{3})$ per iteration or overall for some variants).
While ReX incorporates steps that are computationally demanding, particularly SHAP value estimation, its design aims to harness the expressive power of machine learning models and the insights from XAI to uncover complex causal relationships. This represents a trade-off where enhanced accuracy or the ability to handle non-linearities (as demonstrated in Section 4) may come at a higher computational cost compared to methods with more restrictive assumptions or less exhaustive feature analysis.